OpenStudy (anonymous):

@micahm write the interval as an innequality. Then graph the solutions. [-4,8]

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (anonymous):

you know the difference, and if not I will tell you it. \({\color{black}{ {[a,b]} }}\) is an interval notation for "from a to b", and both a and b are included. \({\color{black}{ {(a,b)} }}\) is an interval notation "from a to b" as well, but, both a and b are NOT included. If I said: \({\color{black}{ {[a,b)} }}\), then it is from a to b, including a and not including b. \({\color{black}{ {(a,b]} }}\), then it is from a to b, not including a and including b. Basic premise: \({\color{black}{ {[} }}\) and \({\color{black}{ {]} }}\) include the number, \({\color{black}{ {(} }}\) and \({\color{black}{ {)} }}\) don't include the number.

OpenStudy (anonymous):

I dont know it.

OpenStudy (anonymous):

so read what I say, and while I you do this, I will out up some examples that can be helpful

OpenStudy (anonymous):

okay so -4 and 8 will both be used in the equation because [a,b] was the given information?

OpenStudy (anonymous):

Look at my first (big) post explaining the interval notations and based on this go through my examples: If my solution was: \({\color{black}{ {4\ge x\ge5} }}\) then the interval notation is, \({\color{black}{ {[4,5]} }}\) If my solution was: \({\color{black}{ {3>x>6} }}\) then the interval notation is, \({\color{black}{ {(3,6)} }}\) If my solution was: \({\color{black}{ {0>x\ge10} }}\) then the interval notation is, \({\color{black}{ {(0,10]} }}\) If my solution was: \({\color{black}{ {7\ge x>11} }}\) then the interval notation is, \({\color{black}{ {[7,11)} }}\)

OpenStudy (anonymous):

when you are done going through thell me if they make sense

OpenStudy (anonymous):

and, I don't have any clue about the given information, since you haven't posted any formation for me to see.

OpenStudy (anonymous):

OHHH okay so my answer would be: \[-4\ge x \ge 8\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

good job: \({\color{black}{ {[-4,8]} }}\) interval, \({\color{black}{ {\Longrightarrow} }}\) \({\color{black}{ {-4\ge x \ge 8} }}\) inequility.

OpenStudy (anonymous):

it should be: \({\color{black}{ {-4\le x \le 8} }}\)

OpenStudy (anonymous):

my error, must admit

OpenStudy (anonymous):

because x is greater than 4, and less than 8,

OpenStudy (anonymous):

and NOT less than 4 and greater than 8